help find the fourth term of the geometric sequence whose first term is 7 and whose common ration is 5
a, =

Answer by mathstutor458(57)

first term,a = 7
common ration , r=15
nth term formula = a1*r^(n-1)
4th term = 7*(15^(4-1))= 7* 15^3 = 23625 -->answer